### v1

# Dimensional crossover of class D real-space topological invariants

Abstract
The topological properties of a material depend on its symmetries, parameters, and spatial dimension. Changes in these properties due to parameter and symmetry variations can be understood by computing the corresponding topological invariant. Since topological invariants are typically defined for a fixed spatial dimension, there is no existing framework to understand the effects of changing spatial dimensions via invariants. Here, we introduce a framework to study topological phase transitions as a system’s dimensionality is altered using real-space topological markers. Specifically, we consider Shiba lattices, which are class D materials formed by magnetic atoms on the surface of a conventional superconductor, and characterize the evolution of their topology when an initial circular island is deformed into a chain. We also provide a measure of the corresponding protection against disorder. Our framework is generalizable to any symmetry class and spatial dimension, potentially guiding the design of materials by identifying, for example, the minimum thickness of a slab required to maintain three-dimensional topological properties.
I Introduction
The possible topological properties of non-interacting Hamiltonians are dictated by their spatial dimension, and discrete and crystal symmetries [1]. In particular, changing the dimension of a system while keeping its discrete symmetries unaltered generally yields a topological phase transition. Examples in condensed matter systems include transitions from a trivial insulator to a quantum spin Hall (QSH) phase in HgTe quantum wells [2, 3], and transitions from three-dimensional topological insulators with two-dimensional surface states to QSH states with one-dimensional helical states [4, 5, 6, 7].
Such topological transitions due to dimensional shifts have previously been characterized theoretically by analyzing edge states in finite-size systems, the topological invariants based on band representation [8, 9], and transport properties [10]. However, a real-space framework capable of illuminating how a system’s topology changes during a dimensional shift has traditionally been missing due to the dearth of local topological markers needed to probe the topology of systems lacking translational symmetry. Such a theory could be useful to guide the design of materials and next-generation devices, for example, to identify how thin a slab can be while maintaining its three-dimensional topological properties, or how thick a nanowire can be while maintaining its edge Majorana modes.
Figure 1: (a) Magnetic adatoms (filled circles) arranged in a circular island (a) and a chain (b) geometry, embedded in a two-dimensional superconductor whose sites are represented by empty circles. Panels (c) and (d) show the zero-energy local density of states (LDOS) for the circular Shiba island and chain, respectively, showing a Majorana mode at their boundary. Panels (e) and (f) show the LDOS as a function of energy at a bulk site (dashed black line) and at a site on the island’s and chain’s edges (solid orange line). The latter displays zero-energy peaks. The parameters used for the island and chain are $N=20$, $\mu=-4$, $\alpha=0.45$, $\Delta=2.0$, and $JS=4.8$
In this paper, we introduce a scheme to study topological phase transitions induced by the change in a system’s spatial dimensions via local topological markers. We will focus on the symmetry class D in the Atland-Zirnbauer classification [11]. Hamiltonians in this class have particle-hole symmetry, $\mathcal{C}$, and broken time-reversal and chiral symmetries. In one dimension (1D), the topological invariant corresponds to $\mathcal{Z}_{2}$, while in two dimensions (2D) to $\mathcal{Z}$ [12, 13]. Our aim is to study the evolution of the $\mathcal{Z}$ invariant into $\mathcal{Z}_{2}$ as one of the dimensions of the system is reduced. Some topological superconductors are examples of class D systems [14].
In particular, we consider magnet-superconductor hybrid systems, Shiba lattices, consisting of time-reversal-symmetry-breaking magnetic atoms placed on the surface of a superconductor arranged as 1D chains [15, 16] or in 2D lattices [17, 18]. Shiba lattices can be assembled with high precision in the laboratory and have shown signatures of topological superconductivity. 1D Shiba chains have been fabricated using ferromagnetic transition metal atoms (such as iron, chromium, manganese, and cobalt) placed on the surface of lead, rhenium, niobium or $\beta$-Bi2Pd [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In 2D, Shiba lattices have been realized using iron adatoms on rhenium [30], flakes of chromium tribromide placed on niobium diselenide [31, 32], and cobalt-silicon islands on Si(111) covered by lead [33]. Additionally, arrays of chromium atoms on niobium have shown signatures of mirror symmetry-protected topological superconductivity [34].
The task of tracking the topological properties of the Shiba lattice across spatial dimensions is well suited to the spectral localizer [35, 36, 37], a real-space numeric K-theory approach that characterizes the local topological properties of finite-size systems, without requiring any notion of lattice periodicity. The spectral localizer approach has been employed to identify topological phenomena in gapless systems [38, 39, 40, 41], nonlinear materials [42], non-Hermitian systems [43, 44], photonic systems [45, 46, 47], quasicrystals [48, 49, 50], time-depedent systems [51, 52], disordered superconductors [53] and heterostructures [54].
We introduce two local topological markers derived from the spectral localizer, a Chern number, and a $\mathcal{Z}_{2}$ invariant that characterizes two and one-dimensional Shiba lattices embedded in two-dimensional superconductors. As these markers do not rely on projecting to an occupied subspace, they are endowed with a measure of the corresponding local topological protection related to the local spectral gap. The local nature of the markers permits us to track their changes and corresponding local protection as the dimension of the Shiba lattice changes from two to one dimension. The real-space topology in 2D Shiba lattices has been characterized using a global position-space invariant [55, 56], and $\mathcal{Z}_{2}$ local markers for nonchiral phases in odd dimensions have been recently introduced [57]. Still, so far, there is no local marker to diagnose the topology of 1D Shiba lattices embedded in a two-dimensional superconductor.
We will next introduce the specific model we consider, define the local topological markers, and present our results. We conclude this paper with a discussion of possible extensions of our work to other physical systems and other symmetry classes.
II Shiba lattice model Hamiltonian
A general mean-field tight-binding Hamiltonian for a Shiba lattice on a two-dimensional superconductor is given by [18, 56]
$\mathcal{H}$ $=-t\sum_{\mathbf{r},\mathbf{r}^{\prime},\sigma}c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}^{\prime}\sigma}-\mu\sum_{\mathbf{r},\sigma}c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}\sigma}+\Delta\sum_{\mathbf{r}}c^{\dagger}_{\mathbf{r}\uparrow}c^{\dagger}_{\mathbf{r}\downarrow}+h.c.$
$+i\alpha\sum_{\mathbf{r},\bm{\delta},a,b}c^{\dagger}_{\mathbf{r}a}\left(\bm{\delta}\times\bm{\sigma}\right)^{z}_{ab}c_{\mathbf{r}+\bm{\delta}b}+J\sum_{\mathbf{R},a,b}\mathbf{S}\cdot c^{\dagger}_{\mathbf{R}a}\bm{\sigma}_{ab}c_{\mathbf{R}b},$ (1)
where $c^{\dagger}_{\mathbf{r}\sigma}$ creates an electron at site $\mathbf{r}$ with spin $\sigma$, $\bm{\sigma}=\{\sigma_{x},\sigma_{y},\sigma_{z}\}$ are the Pauli matrices, $t$ is the hopping amplitude between nearest-neighbor lattice sites $\bm{\delta}$, $l=|\bm{\delta}|$, $\mu$ is the chemical potential, $\alpha$ is the spin-orbit coupling strength, and $\Delta$ is the superconducting s-wave pairing. The magnetic adatoms are located at positions $\mathbf{R}$ and are described as classical magnetic moments $\mathbf{S}$ aligned perpendicular to the superconductor lattice plane and coupled to the electron spin via the exchange coupling $J$.
In Fig. 1, we consider two magnetic adatom geometries: a circular Shiba island (a) and a chain (b), both embedded in a two-dimensional (square lattice) superconductor. We compute the local density of states (LDOS) at site $i$ and energy $E$,
$\rho_{i}(E)=-\frac{1}{\pi}\sum_{\sigma}\textrm{Im}\left[G^{\sigma\sigma}_{ii}(E)\right]$ (2)
using the Chebyshev–Bogoliubov–de Gennes method [58], where $G^{\sigma\sigma}_{ii}(E)$ is the Green’s function. Panels (c) and (d) show the zero-energy LDOS, which shows Majorana modes localized at corresponding 1D and 0D edges for the circular island and chain, respectively. In panels (e) and (f), we show the LDOS as a function of energy at a bulk site and at an edge site of the magnetic adatom geometries. The former case shows the superconducting gap, while the latter shows the zero-energy peak. We considered here the representative parameters $N=20$, where $N$ is the size of the superconductor’s lattice, $\mu=-4$, $\alpha=0.45$, $\Delta=2$, and $JS=4.8$.
Previously, the real-space topological properties of finite-size two-dimensional Shiba lattices have been characterized using real-space Chern numbers [55], including hybrid structures composed of networks of chains attached to islands [56]. Topological quantum numbers for one-dimensional chains based on the scattering matrix have been introduced [15]. A real-space topological invariant for one-dimensional time-quasiperiodic Majorana modes has been recently discussed in Ref. [52]. In contrast, the local topological marker we leverage here applies to one-dimensional Shiba chains embedded in a two-dimensional superconductor, such as the case in Fig. 1(d), relevant for experiments, and is endowed with a local measure of the topological protection.
In the next section, we will introduce the local topological marker that permits us to diagnose the topology of finite-size islands and chains of magnetic adatoms embedded in a two-dimensional superconductor and the crossover from 2D to 1D topology.
III Spectral localizer
In this section, we introduce the local topological markers for 2D and 1D class D Hamiltonians. For a detailed tutorial on the spectral localizer and Clifford pseudospectrum (in the context of photonic systems), see Ref. [59]. Broadly, the spectral localizer framework seeks to classify material topology by first translating a system’s description into a single Hermitian matrix and then diagnose its topology via matrix homotopy, i.e., the study of which types of invertible matrices can be path-connected such that every element in the path shares the same type (e.g., Hermiticity) and remains invertible. As such, the spectral localizer framework does not rely on vector bundles, though it is provably equivalent in some cases [36].
For a d-dimensional system with open boundaries described by the Hamiltonian $\mathcal{H}$ defined on a finite-dimensional Hilbert space, the spectral localizer is defined as [35]
$L_{\bm{\lambda}}$ $(\bm{X},\mathcal{H})=\sum_{j=1}^{d}\kappa(X_{j}-x_{j}\mathcal{I})\otimes\Gamma_{j}+\left(\mathcal{H}-E\mathcal{I}\right)\otimes\Gamma_{d+1},$ (3)
where the matrices $\Gamma_{i}$ generate a Clifford representation, $\mathcal{I}$ is the identity matrix, $X_{j}$ are the position operators, and $\kappa>0$ is a scaling parameter with units of energy divided by length. Intuitively, $\kappa$ is chosen to balance the position and energy contributions to the spectrum of $L_{\bm{\lambda}}$. For bounded Hamiltonians, such as those describing tight-binding models, a proven range of validity exists [37]. Generally, for systems with gapped bulk spectra, this scaling parameter can be set by the bulk gap width $E_{\textrm{gap}}$ and the smallest length of the finite system $L$ as $\kappa\sim E_{\textrm{gap}}/L$. In comparison with other theories of local topological markers [60, 61], setting $\kappa$ is similar to the choice of integration radius used by those markers. Finally, $\bm{\lambda}=(\bm{x},E)$ is the point in position-energy space at which we are probing the local topological properties.
The spectral localizer $L_{\bm{\lambda}}(\bm{X},\mathcal{H})$ encodes both the local topological invariant and a measure of its local protection. In particular, as this framework’s local markers are all rooted in matrix homotopy arguments, the spectral localizer locally quantifies topological protection by how close a system’s $L_{\bm{\lambda}}(\bm{X},\mathcal{H})$ is to being non-invertible at a chosen $\bm{\lambda}$. In other words, the local gap
$\mu_{\bm{\lambda}}(\bm{X},\mathcal{H})=\text{min}\left\{\left|\Sigma\left(L_{\bm{\lambda}}(\bm{X},\mathcal{H})\right)\right|\right\},$ (4)
[…]
In the following sections, we specialize the spectral localizer for the 2D and 1D cases, and define in each case the property that characterizes whether the Hamiltonian and position operators of a class D system can be continued to an atomic limit, i.e., a set of commuting matrices, or not, without breaking parity and while maintaining a non-zero local gap.
Figure 2: (a) Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|$ and (b) 2D localizer gap as a function of position for the 2D island geometry shown in Fig. 1(a), for the same parameters. The spectral localizer calculations use $\kappa=0.05(t/l)$.
III.1 2D invariant
For 2D systems, the Pauli matrices can be chosen as the Clifford representation in Eq. (3), yielding
$L^{(2D)}_{\bm{\lambda}=(x,y,E)}$ $(X,Y,\mathcal{H})=\left(\mathcal{H}-E\mathcal{I}\right)\otimes\sigma_{z}$
$+\kappa\left[(X-x\mathcal{I})\otimes\sigma_{x}+(Y-y\mathcal{I})\otimes\sigma_{y}\right].$ (5)
For a Hamiltonian in class D, the local topological marker at the energy-position point $\bm{\lambda}=(x,y,E)$ corresponds to
$C_{\bm{\lambda}}^{\textrm{L}}=\frac{1}{2}\text{sig}\{L^{(2D)}_{\bm{\lambda}}\left(X,Y,\mathcal{H}\right)\},$ (6)
where $\text{sig}\{L^{(2D)}_{\bm{\lambda}}\left(X,Y,\mathcal{H}\right)\}$, corresponds to the signature of the spectral localizer, the difference between the number of positive and negative eigenvalues. When $\mu_{\bm{\lambda}}$ is nonzero, $C_{\bm{\lambda}}^{\textrm{L}}$ takes a well-defined integer value. In Fig. 2, panel (a), we plot $|C_{\bm{\lambda}}^{\textrm{L}}|$ as a function of position for the Shiba lattice in Fig. 1 (a). In panel (b), we show the corresponding local gap, $\mu_{\bm{\lambda}}$.
Figure 3: Topological phase diagram as a function of the chemical potential $\mu$ and $J$, for $N=20$, and $\alpha=0.45$. Panel (a) shows the case $\Delta=1.2$, while panel (b) $\Delta=0.4$. The color code indicates the local topological marker at the center of a circular island, $|C_{\bm{\lambda}}^{\textrm{L}}|$, and at the center of a chain, $z_{2,\bm{\lambda}}$, using the same parameters for both geometries. Topological transitions not captured in the diagram can occur as the geometry changes from 2D to 1D. The spectral localizer calculations use $\kappa=0.05(t/l)$.
Figure 4: (a)-(e): Arrangements of magnetic adatoms interpolating from the two-dimensional circular island to the one-dimensional chain geometries. (f)-(j): 1D (black solid line) and 2D (dashed orange lines) localizer spectrum $\Sigma(L_{\bm{\lambda}})$ as a function of position $x$ for $y=N/2$. (k)-(o): Local Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|$ (orange squares) and 2D localizer gap $\text{min}\{|\Sigma(L^{(2D)}_{\bm{\lambda}})|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. (p)-(t): Local $z_{2,\bm{\lambda}}$ invariant (black circles) and 1D localizer gap $\text{min}\{\left|\Sigma(L^{(1D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. The parameters used are the same as in Fig. (1). The spectral localizer calculations use $\kappa=0.05(t/l)$.
III.2 1D invariant
The 1D spectral localizer can also be constructed using the Pauli matrices as the choice of Clifford representation as
$L^{(1D)}_{\bm{\lambda}=(x,E)}(X,\mathcal{H})=\left(\mathcal{H}-E\mathcal{I}\right)\otimes\sigma_{y}+\kappa(X-x\mathcal{I})\otimes\sigma_{x},$ (7)
where $\mathcal{H}$ is also given by Eq. (1). Note that the choice of $\mathcal{H}$ need not change bewteen using $L^{(2D)}_{\bm{\lambda}}$ and $L^{(1D)}_{\bm{\lambda}}$. Instead, the dimensional reduction is performed through the omission of the second position operator $Y$ in the definition of $L^{(1D)}_{\bm{\lambda}}$. As such, $L^{(1D)}_{\bm{\lambda}}$ can be equally applied to 1D chains as well as effective 1D chains embedded in 2D materials (or any material of arbitrarily higher dimension). Similarly, one can also define $L^{(1D)}_{\bm{\lambda}}$ in terms of other position operators such as $Y$ or $X+Y$. However, for concreteness, we arrange our Shiba chains along the x-direction in the 2D superconductor.
The 1D local topological invariant in class D is defined by first transforming $\mathcal{H}$ to be in a basis in which it is purely imaginary, so that $\mathcal{H}^{\mathsf{T}}=-\mathcal{H}$. Such a basis must exist due to the system’s particle-hole symmetry. In this basis for $\mathcal{H}$, and noting the choice of Pauli matrices used in Eq. (7), the 1D spectral localizer is both block off-diagonal and real symmetric, $(L^{(1D)}_{\bm{\lambda}})^{\dagger}=(L^{(1D)}_{\bm{\lambda}})^{\mathsf{T}}=L^{(1D)}_{\bm{\lambda}}$. Thus, all of its eigenvalues come in pairs, and when $L^{(1D)}_{\bm{\lambda}}$ is invertible, its matrix homotopy can be classified using only one of its off-diagonal blocks (which removes this spectral pairing) as
$z_{2,\bm{\lambda}}=\text{sign}\left[\text{det}\left\{\kappa(X-x\mathcal{I})-i(H-E\mathcal{I})\right\}\right],$ (8)
see Theorem D1 in Ref. [62] for an accessible explanation.
In the next section, we calculate the 1D and 2D topological invariants for a chain and an island, respectively, to define a topological phase diagram. We then characterize the topological transitions induced by the change in dimension, as a Shiba island is deformed into a Shiba chain.
Figure 5: (a)-(e): Arrangements of magnetic adatoms interpolating from the two-dimensional circular island to the one-dimensional chain geometries. (f)-(j): 1D (black solid line) and 2D (dashed orange lines) localizer spectrum $\Sigma(L_{\bm{\lambda}})$ as a function of position $x$ for $y=N/2$. (k)-(o): Local Chern number, $|C_{\bm{\lambda}}^{\textrm{L}}|$, (orange squares) and 2D localizer gap $\text{min}\{\left|\Sigma(L^{(2D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. (p)-(t): Local $z_{2,\bm{\lambda}}$ invariant (black circles) and 1D localizer gap $\text{min}\{\left|\Sigma(L^{(1D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. The parameters are $N=20$, $\mu=1.0$, $\alpha=0.45$, $\Delta=1.2$, and $JS=2.579$. The spectral localizer calculations use $\kappa=0.05(t/l)$.
IV Results
Leveraging the local topological invariants for Class D systems, we explore the Shiba lattice’s parameter space for 1D and 2D geometries and consider values previously used in the literature for this model [63, 55, 64, 56]. To this end, we consider two geometries, a circular island, and a chain, and compute the corresponding local topological invariants (the Chern number $C_{\bm{\lambda}}^{\textrm{L}}$ and $z_{2,\bm{\lambda}}$, respectively) at the system’s center. The results are plotted in Fig. 3. Using the same parameters for both geometries, we find regions where the Shiba island and chain are both topological, regions where the Shiba island is topological and the chain is trivial, regions where the island is trivial and the chain is topological, and regions where both geometries are trivial. Note that topological transitions not captured in the diagram can occur as the geometry changes from 2D to 1D. For example, the transition from $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ to $z_{2,\bm{\lambda}}=-1$ as the circular island is flattened occurs via a change to $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ for a narrow island, as we discuss later.
Next, we consider representative cases in the parameter space and study the topological invariants as the Shiba geometry transitions from two-dimensional to one-dimensional.
IV.1 Topological invariant dimensional crossover
Figure 6: Length of the cut along $x$ for $y=N/2$, as shown in Fig. (4), for which the local topological invariant is nontrivial, and the local gap is larger than $0.02t$, $d^{top}$, divided by the number of sites along the cut, $d^{tot}$. The parameters for panel (a) are the same as in Fig. 4. In panel (b), we show the case Fig. 9 with parameters $N=20$, $\mu=0$, $\alpha=0.8$, $\Delta=1.2$, and $JS=2$. Panel (c) shows the case in Fig. 5, the parameters are $N=20$, $\mu=1.0$, $\alpha=0.45$, $\Delta=1.2$, and $JS=2.579$. The spectral localizer calculations use $\kappa=0.05(t/l)$.
First, we consider the case where both the Shiba island and chain are topological with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ and $z_{2,\bm{\lambda}}=-1$, respectively. In Fig. 4, top row, we plot the set of magnetic adatom geometries considered, which interpolate between a Shiba circular island [(a)] and a Shiba chain [(e)] by reducing the number of magnetic adatoms along the y-direction. In the second row, we plot the spectra of the 1D (black solid lines) and 2D (dashed orange lines) spectral localizers, $\Sigma(L_{\bm{\lambda}}(\bm{X},\mathcal{H}))$, computed as a function of position $x$, for $y=N/2$ and zero energy.
For the Shiba island, panel (f), $\Sigma(L^{(2D)}_{\bm{\lambda}}(\bm{X},\mathcal{H}))$ crosses zero at the edge of the magnetic adatom island, transitioning from trivial $C_{\bm{\lambda}}^{\textrm{L}}=0$ to topological with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ (panel (k)). The finite localizer gap is maximal at the center of the island. On the other hand, the spectrum of the 1D spectral localizer, $\Sigma(L^{(1D)}_{\bm{\lambda}}(\bm{X},\mathcal{H}))$, displays a much smaller local gap at the location of magnetic adatoms. While there are sites where the topological local marker is topological, $z_{2,\bm{\lambda}}=-1$, there is no local gap protection against variations in the Hamiltonian $\delta\mathcal{H}$ or position operators $\delta X$. Therefore, the 1D topology is trivial, and the system is characterized by a two-dimensional topology via the local Chern number, $C_{\bm{\lambda}}^{\textrm{L}}$.
We now reduce the number of magnetic adatoms along the y-direction, flattening the island, as shown in panel (b). Similar trends in the local markers and the localizer gaps are obtained. However, the number of sites with a nontrivial Chern number is reduced compared with the circular island’s case. Additionally, the 2D localizer gap is now smaller. The trend continues for a flatter island, as shown in panel (c). However, for the parameters considered, the three-atom-tall island exhibits a behavior change. The Chern number is now nonzero only at locations where the 2D localizer gap vanishes, panel (n), while there are sites with a nontrivial $z_{2,\bm{\lambda}}$ invariant protected by a finite 1D localizer gap, panel (s). This marks the transition from 2D to 1D topology in the system. For the Shiba chain, panel (e), the $z_{2,\bm{\lambda}}$ invariant changes from trivial to nontrivial at the edge of the chain. As the Shiba island reduces to a chain, $\Sigma(L^{(2D)}_{\bm{\lambda}=(x,y=N/2,E=0)}(\bm{X},\mathcal{H}))$ approaches $\Sigma(L^{(1D)}_{\bm{\lambda}=(x,E=0)}(\bm{X},\mathcal{H}))$.
We now consider a set of parameters that leads to a topological island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into a topological chain. We show our results in Fig. 5. There are two main differences with respect to the previous example we discussed. First, there is a topological transition from $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ to $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ when the Shiba island shrinks from seven to five atoms tall. Second, for the three-atom-tall island, there is a coexistence of 2D and 1D topology with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ and $z_{2,\bm{\lambda}}=-1$, respectively. However, as we will discuss later, when including disorder in the chemical potential, the two-dimensional local gap is more sensitive to disorder than the 1D local gap.
Finally, in Appendix A, we show plots for a topological island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into the trivial chain. In this case, the number of sites with nontrivial topology decreases together with the local gap protection, until it becomes trivial in both 1D and 2D. We present a graphical summary of our calculations in Fig. 6, where we plot the number of sites with a topological local marker, $d^{top}$, divided by the total number of sites on the $x$-axis, $d^{tot}$, both for 1D and 2D, as a function of the island’s width to length ratio. We only consider sites topological if the local gap is larger than $0.02t$. The transition from 2D topology to 1D topology, as characterized by the spectral localizer, is given by the crossing of the traces.
Figure 7: (a) 2D localizer gap for a disordered circular Shiba island (as shown in the inset) along position $x$ and $y=N/2$, with disorder strength $W/t=1.2$. The dots show the average over 100 disorder realizations, while the shaded area corresponds to ±$\pm 1$ standard deviations. Panel (b) shows the corresponding local maker, $|C_{\bm{\lambda}}^{\textrm{L}}|$, which is trivial. (c) 1D localizer gap for a disordered Shiba chain (as shown in the inset) along position $x$ and $y=N/2$, also for $W/t=1.2$, corresponding to the chain’s position. Panel (d) shows the corresponding local marker, $z_{2,\bm{\lambda}}$, which remains topological. The spectral localizer calculations use $\kappa=0.05(t/l)$.
IV.2 Effects of disorder
To test the robustness of our result, we consider disorder in the local chemical potential. The additional term we include in the Hamiltonian is given by
$\mathcal{H}_{dis}=-\sum_{\mathbf{r},\sigma}\delta\mu(\mathbf{r})c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}\sigma},$ (9)
where $\delta\mu(\mathbf{r})$ is taken to be an uncorrelated, uniformly distributed random variable, $\delta\mu(\mathbf{r})\in[-W,W]$. Shiba lattices could also present disorder in the adatom coupling strength, spin orientation, and possibly missing magnetic atoms in a given geometry [55]. In this study, we do not calculate the superconducting order parameter self-consistently in the presence of disorder [65]. In Fig. 7, we consider the disorder strength $W/t=1.2$ and compute the topological markers and local gaps for an island and a chain geometry. We average over a set of 100 disorder realizations. Panels (a) and (c) show the 2D and 1D local gaps, respectively. The dots show the average while the shaded area shows ±$\pm 1$ standard deviations. For this disorder strength, the 2D local gap is closed along the position $x$ path considered. The system is therefore 2D-trivial. However, the 1D local gap remains open, and system remains 1D-topological, as shown in panel (d), which shows $z_{2,\bm{\lambda}}$.
Now, we consider the effect of disorder in the case where 1D and 2D topologies coexist in the clean case, as shown in Fig. 5, panels (n) and (s). We consider a disordered three-atom-tall Shiba island with disorder strength $W/t=0.3$ and average over 100 disorder realizations. The results are shown in Fig. 8. Panel (a) shows the 2D localizer gap. The solid gray line corresponds to the clean case, the blue dots to the average localizer gap, and the shaded area indicates ±$\pm 1$ standard deviations. The corresponding local Chern number, $|C_{\bm{\lambda}}^{\textrm{L}}|$, is shown in (b). The small gap protecting the topology in the clean limit is now significantly reduced. Panel (c) shows the 1D localizer gap, which remains open across the region with nontrivial $z_{2,\bm{\lambda}}=-1$. Therefore, the chemical potential disorder favors the 1D topology.
In the next section, we present the conclusions of our study and discuss possible extensions of our work to other physical systems and other symmetry classes.
Figure 8: (a) 2D localizer gap for a disordered three-atom-tall Shiba island, as shown in the inset, with along position $x$ and $y=N/2$, with disorder strength $W/t=1.2$. The dots show the average over 100 disorder realizations, while the shaded area corresponds to ±$\pm 1$ standard deviations. Panel (b) shows the corresponding local maker, $|C_{\bm{\lambda}}^{\textrm{L}}|$, with orange squares. The clean limit results are shown in gray. (c) 1D localizer gap for a disordered Shiba chain along position $x$ and $y=N/2$, also for $W/t=1.2$, corresponding to the chain’s position. Panel (d) shows the corresponding local marker, $z_{2,\bm{\lambda}}$, which remains topological. The gray lines correspond to the clean limit. The spectral localizer calculations use $\kappa=0.05(t/l)$.
V Conclusions and outlook
We have established a framework based on the spectral localizer, a real-space numeric K-theory approach, to characterize topological transitions due to spatial dimension changes. In particular, we introduced local topological markers for systems in class D, with particle-hole symmetry and broken time-reversal and chiral symmetry, and applied them to Shiba lattices as experimentally-relevant examples. We considered a 2D circular Shiba island and deformed it into a 1D chain by reducing its size along one of its dimensions. By computing the local markers across the dimension change, we found that the transition from 2D topology to 1D topology occurs via a reduction in the number of sites with Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|>0$ together with a decrease in the localizer gap protecting it. At a critical thickness, we found the emergence of nontrivial 1D topology, marked by $z_{2,\bm{\lambda}}=-1$, protected by a nonzero 1D localizer gap whose size increases for the 1D chain geometry. We tested the robustness of our results against disorder in the chemical potential. We found that the 1D localizer gap, and thus the 1D topology, is more robust than its 2D counterpart.
The framework introduced can be readily generalized to other symmetry classes. For example, the spectral localizer, Eq. (3), can be defined for a 3D topological insulator and a 2D quantum spin Hall system. The corresponding topological invariants would be a Chern number, analogous to Eq. (6), and the spin Chern number. More generally, the framework could help guide the design of topological materials for device applications. One of the strengths of the spectral localizer framework is its numerical efficiency, which enables it to be applied directly to experimental systems without developing a low-energy approximation [66]. As such, our results may be useful for the design and analysis of topological superconducting nanowires [67].
Acknowledgements.
We thank Annica M Black-Schaffer and Arnob Kumar Ghosh for insightful discussions. A.C. acknowledges support from the Laboratory Directed Research and Development program at Sandia National Laboratories. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology &\&& Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. Research by T.L. was supported in part by the National Science Foundation, award DMS-2349959. Research by T.L. was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-25-1-0052. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government.
Appendix A Appendix A
In this appendix, we consider the case of a Shiba island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into a trivial chain, with $z_{2,\bm{\lambda}}=1$. The topological dimensional crossover occurs for the three-atoms-tall island, analogous to the cases discussed in the main text, but in this case, the system is trivial.
Figure 9: (a)-(e): Arrangement of magnetic adatoms interpolating from the two-dimensional circular island to the one-dimensional chain. (f)-(j): 1D (black solid line) and 2D (dashed orange lines) localizer spectrum $\Sigma(L_{\bm{\lambda}})$ as a function of position $x$ for $y=N/2$. (k)-(o): Local Chern number, $|C_{\bm{\lambda}}^{\textrm{L}}|$, (orange squares) and 2D localizer gap $\text{min}\{\left|\Sigma(L^{(2D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. (p)-(t): Local $z_{2,\bm{\lambda}}$ invariant (black circles) and 1D localizer gap $\text{min}\{\left|\Sigma(L^{(1D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. The parameters used are $N=20$, $\mu=0$, $\alpha=0.8$, $\Delta=1.2$, and $JS=2$. The spectral localizer calculations use $\kappa=0.05(t/l)$.